A van travels 240 miles on 12 gallons of gas. Write and solve a proportion to find how many gallons the van needs to travel 420 miles.
Answer by ewatrrr(24785)


Hi
Let x represent gallons the van needs to travel 420 miles
Question Sets Up the Proportion
240/12 = 420/x |Cross Multiplying
240x = 12*420
x = 12*420/240
x = 21 gal, the amount the van needs to travel 420 miles
CHECKING our Answer***
240/12 = 420/21 = 20/1